Under vector addition,

α+(β+γ)=(β+α++γ is called the ____ law.


1
Expert's answer
2021-06-29T04:41:54-0400

Solution:

Take α+(β+γ)

=(α+β)+γ [Associative property of addition]

=(β+α)+γ [Commutative property of addition]

Thus, Under vector addition,

α+(β+γ)=(β+α)+γ is called the associative and commutative of addition law